OpenStudy (anonymous):

I'll fan. Just help me with this: From the top of a 40m tower, an observer measures the angle of depression to a boat in the water to be 28 degrees. How far is the boat from the base of the tower?

OpenStudy (aaronandyson):

First,Draw picture.

OpenStudy (aaronandyson):



OpenStudy (aaronandyson):

62 = 90 - 28 Properties of triangle.

OpenStudy (aaronandyson):

Tan 62 = x/40 Where x is opposite side 40 is adjacent side

OpenStudy (aaronandyson):

Value of tan 62 is 1.88

OpenStudy (aaronandyson):

1.88 = x/40

OpenStudy (aaronandyson):

Moving to 40 to the LHS

OpenStudy (aaronandyson):

1.88*40 = x

OpenStudy (aaronandyson):

18.8*4 = x I remove the zero of forty and gave it to 1.88 which is now 18.8